OpenStudy (anonymous):

I need help understanding how we get the numerator for this question I understand the mixed fraction part but not the numerator 2 /3+1/4

OpenStudy (helder_edwin):

the easiest (but not elegant at all) way to do addition of fractions is this \[ \large \frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd} \] after you do this, just simplify